# Two-Timescale Gradient Descent Ascent Algorithms for Nonconvex Minimax Optimization

### v1

@@ 6 Applications and Empirical Results
We discuss our evaluation of TTGDA and TTSGDA in two domains – robust regression and Wasserstein generative adversarial networks (WGANs). All of algorithms were implemented on a MacBook Pro with an Intel Core i9 2.4GHz and 16GB memory.
@@ 6.1 Robust logistic regression
We will consider the problem of robust logistic regression with nonconvex penalty functions. Indeed, given a set of samples $\{(\mathbf{a}_{i},b_{i})\}_{i=1}^{N}$ where $\mathbf{a}_{i}\in\mathbb{R}^{d}$ is the feature of the $i^{\textnormal{th}}$ sample and $b_{i}\in\{1,-1\}$ is the label, the problem can be written in the form of Eq. (1.1) with
@@ 6.2 WGANs with linear generators
$f(\mathbf{x},\mathbf{y})=\mathbb{E}_{(a^{\textnormal{real}},z)}\left[D_{\mathbf{y}}(a^{\textnormal{real}})-D_{\mathbf{y}}(G_{\mathbf{x}}(z))-\lambda\|\mathbf{y}\|^{2}\right].$
@@ 6.3 WGANs with nonlinear generators
Following the setup of Lei et al. (2020) and Yang et al. (2022b), we will consider the WGAN with a ReLU neural network as generator. For ease of comparison, we leave all the problem settings identical to last paragraph and view $\mathbf{x}$ as the parameters of a small neural network (1 hidden layer with 5 neurons and ReLU activations).
Figure 3 shows that TTSGDA is worse with Adam and RMSprop in this complex setting. This makes sense since the nonlinear generators result in a richer structure and the adaptive algorithms are highly optimized for handling such structure. The vanilla implementation of TTSGDA can not achieve the satisfactory performance and can be enhanced by exploiting the advanced tools of regularization and adaptiveness (Yang et al., 2022a, b; Li et al., 2023). For more details on the implementation and empirical performance for TTSGDA in training GANs and its variants, we refer to Heusel et al. (2017).
@@ 7 Conclusion
We proposed two-timescale GDA and SGDA algorithms (TTGDA and TTSGDA for short) for solving a class of nonconvex minimax optimization problems. The complexity bounds were established for TTGDA and TTSGDA in terms of the number of (stochastic) gradient evaluations, demonstrating the efficiency of these two algorithms in theory. Our theoretical analysis is of interest to both optimization and machine learning communities as it not only provides a new combination of a two-timescale update rule and GDA/SGDA but supplies a novel proof technique that applies to concave optimization problems with a slowly changing objective functions. Future directions include the investigation of lower bounds for gradient-based algorithms in the nonconvex-concave settings and the extension of upper bounds for TTGDA and TTSGDA in the structured nonconvex-nonconcave settings.
@@ Appendix A Proofs for Technical Lemmas
To further facilitate the readers, we give a brief description of the extragradient algorithm. It was introduced by Korpelevich [1976] and has been recognized as one of the most popular methods for solving the minimax optimization problems. In the convex-concave setting, its global convergence rate guarantees were established for both averaged iterates [Nemirovski, 2004, Mokhtari et al., 2020a, b] and last iterates [Gorbunov et al., 2022, Cai et al., 2022]. When applied to solve the optimization problem of $\min_{\mathbf{x}}\max_{\mathbf{y}\in\mathcal{Y}}f(\mathbf{x},\mathbf{y})+\ell\|\mathbf{x}-\hat{\mathbf{x}}\|^{2}$, one iteration of the extragradient algorithm is summarized as follows,
@@
Since $\mathbf{y}^{\star}(\cdot)$ is a well-defined function and $\mathcal{Y}$ is a convex and bounded set, the celebrated Danskin’s theorem (see Danskin [1967] and Bertsekas [1971, Proposition A.22]) implies that $\nabla\Phi(\cdot)=\nabla_{\mathbf{x}}f(\cdot,\mathbf{y}^{\star}(\cdot))$. Since $\mathbf{y}^{\star}(\cdot)$ is $\kappa$-Lipschitz and $\kappa\geq 1$, we have
@@ Appendix B Smooth and Nonconvex-Strongly-Concave Setting
Using the Young’s inequality yields
@@
Using the Young’s inequality and that the function $\mathbf{y}^{\star}(\cdot)$ is $\kappa$-Lipschitz, we have
@@
We then prove the results for Algorithm 2. By applying the same argument as used for proving Eq. (B.1), we have
@@ Appendix D Nonsmooth and Nonconvex-Strongly-Concave Setting
Since the function $f(\cdot,\mathbf{y})$ is $L$-Lipschitz for each $\mathbf{y}\in\mathcal{Y}$, we have $\|\mathbf{g}\|\leq L$ for all $\mathbf{g}\in\partial_{\mathbf{x}}f(\mathbf{x},\mathbf{y})$ and each pair of $\mathbf{x}$ and $\mathbf{y}\in\mathcal{Y}$. Applying the same argument as used for proving Lemma C.2 yields
@@
Since the function $f(\mathbf{x}_{t-1},\cdot)$ is $L$-Lipschitz, we have $\|\textbf{g}_{\mathbf{y}}^{t-1}\|\leq L$. Combining this fact with the Young’s inequality yields
@@
Using the Young’s inequality, we have
@@ Appendix E Nonsmooth and Nonconvex-Concave Setting
We then prove the results for Algorithm 2. By applying the same argument as used for analyzing Algorithm 1 but with the second inequalities in Lemmas D.1 and E.2, we have

### v2

@@ 6 Applications and Empirical Results
We discuss our evaluation of TTGDA and TTSGDA in two domains—robust regression and Wasserstein generative adversarial networks (WGANs). All of algorithms were implemented on a MacBook Pro with an Intel Core i9 2.4GHz and 16GB memory.
@@ 6.1 Robust logistic regression
We will consider the problem of robust logistic regression with nonconvex penalty functions. Given a set of samples $\{(\mathbf{a}_{i},b_{i})\}_{i=1}^{N}$ where $\mathbf{a}_{i}\in\mathbb{R}^{d}$ is the feature of the $i^{\textnormal{th}}$ sample and $b_{i}\in\{1,-1\}$ is the label, the problem can be written in the form of Eq. (1.1) with
@@ 6.2 WGANs with linear generators
$f(\mathbf{x},\mathbf{y})=\mathbb{E}_{(a^{\textnormal{real}},z)}\left[D_{\mathbf{y}}(a^{\textnormal{real}})-D_{\mathbf{y}}(G_{\mathbf{x}}(z))-\lambda\|\mathbf{y}\|^{2}\right],$
@@ 6.3 WGANs with nonlinear generators
Following the setup of Lei et al. (2020) and Yang et al. (2022b), we consider the WGAN with a ReLU neural network as generator. For ease of comparison, we retain all the problem settings as in the previous paragraph and view $\mathbf{x}$ as the parameters of a small neural network (one hidden layer with five neurons and ReLU activations).
Figure 3 shows that TTSGDA is worse with Adam and RMSprop in this complex setting. This makes sense since the nonlinear generators result in a richer structure and the adaptive algorithms are highly optimized for handling such structure. The vanilla implementation of TTSGDA does not achieve a satisfactory level of performance and can be enhanced by exploiting the advanced tools of regularization and adaptiveness (Yang et al., 2022a, b; Li et al., 2023). For more details on the implementation heuristics and empirical performance for TTSGDA in training GANs and its variants, we refer to Heusel et al. (2017) for the details.
@@ 7 Conclusion
We proposed two-timescale GDA and SGDA algorithms (TTGDA and TTSGDA for short) for solving a class of nonconvex minimax optimization problems. We established complexity bounds for TTGDA and TTSGDA in terms of the number of (stochastic) gradient evaluations, demonstrating the efficiency of these two algorithms in theory. Our theoretical analysis is of interest to both optimization and machine learning communities as it not only provides a new combination of a two-timescale update rule and GDA/SGDA but supplies a novel proof technique that applies to concave optimization problems with slowly changing objective functions. Future directions of research include the investigation of lower bounds for gradient-based algorithms in the nonconvex-concave settings and the extension of upper bounds for TTGDA and TTSGDA in the structured nonconvex-nonconcave settings.
@@ Appendix A Proofs for Technical Lemmas
To further facilitate the reader’s understanding, we give a brief description of the extragradient algorithm. It was introduced by Korpelevich [1976] and has been recognized as one of the most useful methods for solving minimax optimization problems. In the convex-concave setting, its global convergence rate guarantees were established for both averaged iterates [Nemirovski, 2004, Mokhtari et al., 2020a, b] and last iterates [Gorbunov et al., 2022, Cai et al., 2022]. When applied to solve the optimization problem of $\min_{\mathbf{x}}\max_{\mathbf{y}\in\mathcal{Y}}f(\mathbf{x},\mathbf{y})+\ell\|\mathbf{x}-\hat{\mathbf{x}}\|^{2}$, one iteration of the extragradient algorithm is summarized as follows,
@@
Since $\mathbf{y}^{\star}(\cdot)$ is a well-defined function and $\mathcal{Y}$ is a convex and bounded set, the celebrated Danskin theorem (see Danskin [1967] and Bertsekas [1971, Proposition A.22]) implies that $\nabla\Phi(\cdot)=\nabla_{\mathbf{x}}f(\cdot,\mathbf{y}^{\star}(\cdot))$. Since $\mathbf{y}^{\star}(\cdot)$ is $\kappa$-Lipschitz and $\kappa\geq 1$, we have
@@ Appendix B Smooth and Nonconvex-Strongly-Concave Setting
Using Young’s inequality yields
@@
Using Young’s inequality and the fact that the function $\mathbf{y}^{\star}(\cdot)$ is $\kappa$-Lipschitz, we have
@@
We now prove the results for Algorithm 2. By applying the same argument as used for proving Eq. (B.1), we have
@@ Appendix D Nonsmooth and Nonconvex-Strongly-Concave Setting
Since $f(\cdot,\mathbf{y})$ is $L$-Lipschitz for each $\mathbf{y}\in\mathcal{Y}$, we have $\|\mathbf{g}\|\leq L$ for all $\mathbf{g}\in\partial_{\mathbf{x}}f(\mathbf{x},\mathbf{y})$ and each pair of $\mathbf{x}$ and $\mathbf{y}\in\mathcal{Y}$. Applying the same argument as used for proving Lemma C.2 yields
@@
Since the function $f(\mathbf{x}_{t-1},\cdot)$ is $L$-Lipschitz, we have $\|\textbf{g}_{\mathbf{y}}^{t-1}\|\leq L$. Combining this fact with Young’s inequality yields
@@
Using Young’s inequality, we have
@@ Appendix E Nonsmooth and Nonconvex-Concave Setting
We now prove the results for Algorithm 2. By applying the same argument as used for analyzing Algorithm 1 but with the second inequalities in Lemmas D.1 and E.2, we have